# Fast Polynomial Root-Finding and Approximation of the Eigenvalues of a Matrix

### v36

@@ 18.2 Bit operation cost of our e/i tests and root-finders
In Part IV we decrease the number of subdivision steps of our root-finders from order of $b$ to $O(\log(b))$, thus decreasing bounds (18.3) and (18.4) by a factor of $b$, and arrive at our Main Theorems 2, 3, and 4.
PART IV: Compression of a Disc and Acceleration of Subdivision Root-Finders
19 Overview and two alternative approaches
Classical subdivision iterations converge to the roots with linear convergence rate (see Sec. 2), which is slow when many consecutive iterations process a component $C$ made up of a single suspect square or up to four of them sharing a vertex. In this case the algorithms of [128, 101] dramatically accelerate convergence by means of applying Newton’s or Schröder’s iterations (see Fig. 8), which converge with superlinear rate to the convex hull of all roots lying in $C$. The diameter of the convex hull, however, can be too large, and then [128, 101] apply subdivision iterations again.
The resulting root-finder recursively intertwines subdivision iterations with Newton’s or Schröder’s, and this complicates the complexity estimates for a root-finder. [128, 101] have only estimated the number of ops but not of the bit operations involved. [23, 24] estimated the bit operation complexity, but only under a very strong assumption about the output (see Remark 4.2 and Sec. 12.4).
Figure 8: Five blue marked centers of suspect squares converge to a green-marked root cluster with linear rate; two red marked Schröder’s iterates converge with quadratic rate.
Instead of Newton’s or Schröder’s iterations, one can apply the algorithm of [134], which splits $p(x)$ into the product of two polynomials $f(x)$ and $g(x)$ whose zero sets lie in and outside the minimal disc covering the component $C$, respectively. Then one can apply a root-finder to the polynomials $f=f(x)$ and $g=g(x)$. Splitting of $p(x)$ into factors does not preserve sparsity of $p(x)$, and the factors may have significantly larger coefficient length than $p(x)$. (E.g., split $p(x)=x^{1000}+1$ into two factors of degrees 300 and 700.)
Overall asymptotic bit-complexity for high precision root-finding can still be near-optimal if a root-finder recursively splits $p(x)$ into factors of comparable degrees. This is achieved in [96, 103], but the resulting root-finder is much involved and extensively uses the coefficients of $p$.
The problems caused by splitting disappear if one of the factors, say, $g$ has a small degree; then we can deflate it, approximate its zeros readily, and approximate the zeros of $f(x)=p(x)/g(x)$ by applying to it a black box root-finder and using implicit deflation based on (3.3) or the equation NIR$(p)=$NIR(f)+limit-$(f)+$NIR$(g)$ for $p=fg$.
One would lose optimality, however, by always deflating low-degree factors: to deflate $k$ times one would need at least order of $bkmd$ bit operations overall, that is, would exceed the lower bound $0.25(d+1)db$ of Cor. 24.2 by at least a factor of $k$.
We stay with subdivision framework but instead of Newton’s or Schröder’s iterations, used in [128, 101], apply our novel algorithm that compresses a component isolated from its external roots into a near-minimal disc that keeps all internal roots. This enables us to decrease the estimates of Secs. 12.3 and 18.2 for the overall complexity of root-finding by a factor of $b$, thus completing proof of our Main Theorems 2-4.
We organize our presentation in this part as follows. In the next section we study a tree representing recursive partition of the union of suspect squares in a subdivision process into connected components. In Sec. 21 we devise and analyze our first compression algorithm. In particular we observe in Sec. 21.4 that the whole subdivision process involves at most $12m-$ 1 compression steps, reduced to $O(m)$ e/i tests. In Sec. 22 we prove that in addition it also involves $O(m)$ e/i steps. We reach a near-optimal bound on the bit operation complexity of our root-finders in Sec. 23 by devising and analyzing our compression algorithm and recalling our estimates for the complexity of an e/i step. In that section we also present a fast root-counter in an isolated disc.
We use some results of Part I, Alg. 9.1 of Part II, and Sec. 13 (Background) of Part III. Otherwise our presentation in this part of the paper is self-contained.
20 Component tree and compact components
20.1 Component tree, edge length, non-adjacent suspect squares, and compact components
Recall the definition of the component tree $\mathbb{T}$ of Sec. 4, which represent partition of the union of suspect squares at every subdivision step and representation of such partition with the component tree $\mathbb{T}$ of Sec. 4. The tree has at most $m$ components-leaves that represent the output $\epsilon$-discs and at most $m-1$ other vertices, which represent components-parents, each broken by a subdivision step into at least two components-children. The initial suspect square is the root of the tree $\mathbb{T}$. The edges of the tree connect parents and children, representing subdivision steps between partitions of components. Cor. 20.1 below bounds the number of such steps; one can consider this number the edge length.
Call a pair of suspect squares non-adjacent if they share no vertices.
@@ 20.1 Component tree, edge length, non-adjacent suspect squares, and compact components
Call a component compact if it contains just a single non-adjacent suspect square, that is, consists of at most four suspect squares sharing a vertex (see Fig. 10).
@@
A component is broken in at most $1+\lfloor 0.5+\log_{2}(m)\rfloor$ successive subdivision steps unless it is compact.
20.2 Isolation of compact components
Observation 20.2.
@@ 20.2 Isolation of compact components
21 $\epsilon$-compression: algorithm and analysis
21.1 Compression algorithm: high level description
Recall that the classical subdivision iterations converge slowly where they successively process only compact components (cf. Fig. 8), but then we accelerate this process by interrupting it with compression steps in at most three iterations.
Next we specify compression for a 3.5-isolated unit disc $D=D(0,1)$ covering a component; map (1.3) extends compression to any 3.5-isolated disc $D(c,\rho)$; the map $\rho\epsilon\mapsto\epsilon$ adjusts the error tolerance $\epsilon$ by scaling by $\rho$. Observation 20.2 implies that we can readily ensure $\theta$-isolation at the cost of performing $O(\log(\theta))$ additional subdivision steps.
Algorithm 21.1.
$\epsilon$-compression (high level description).
INPUT: a positive $\epsilon\leq 1$ and a black box polynomial $p(x)$ of a degree $d$ such that [^34] Newton’s and Schröder’s iterations of [128, 101] accelerate convergence of subdivision iterations to a root cluster in a disc if the disc is $5d^{2}$-isolated, while Alg. 21.1 only requires 3.5-isolation (see (21.1)).
$i(D(0,1))\geq 3.5~{}{\rm and}~{}\#(D(0,1))\geq 1.$ (21.1)
@@ 21.1 Compression algorithm: high level description
OUTPUT: an integer $m$, a complex $c^{\prime}$, and a positive $\rho^{\prime}$ such that $m=\#(D(0,1))$, (i) $X(D(c^{\prime},\rho^{\prime}))=X(D(0,1))$, and (ii) $\rho^{\prime}\leq\epsilon$ or $8|z_{g}-z_{h}|>5\rho^{\prime}$ for a pair of roots $z_{g},z_{h}\in D(c^{\prime},\rho^{\prime})$. [Hence the disc $D(c^{\prime},\rho^{\prime})$ is 5/8-rigid.]
INITIALIZATION: Fix $\delta\geq 0$ such that $8\delta\leq\epsilon\leq 1$.
COMPUTATIONS: Compute $m=\#(D(0,1))$, $c^{\prime}$, and $\rho^{\prime}$ such that
$|c^{\prime}-c|\leq\delta~{}{\rm and}~{}\rho\leq\rho^{\prime}\leq\max\{\epsilon,\rho+2\delta\}$ (21.2)
for $\rho:=r_{d-m+1}(c^{\prime})$ and some $c\in CH(X(D(0,1)))$.
21.2 Correctness of compression algorithm
First prove correctness in a simple case where $\epsilon=\delta=0$, $c^{\prime}=c$, $\rho^{\prime}=\rho$.
In this case $c^{\prime}=c\in CH(X(D(0,1)))\subset D(0,1)$, and so $\rho^{\prime}=\rho=r_{d+1-m}(c^{\prime})\leq 2$. Hence $D(c^{\prime},\rho^{\prime})\subseteq D(0,3)$.
Furthermore, #$\#(D(c^{\prime},\rho^{\prime}))=m$ since $\rho^{\prime}=\rho:=r_{d+1-m}(c^{\prime})$, while #$\#(D(0,1))=m$ by assumption.
Now (21.1) implies that #$\#(D(0,3))=m$, which supports claim (i) for the output of Alg. 21.1.
Next let $|c^{\prime}-z_{g}|=\rho:=r_{d+1-m}(c^{\prime})$ for a root $z_{g}$. Then all $m$ roots in $D(0,1)$ also lie in the disc $D(c^{\prime},\rho)=D(c^{\prime},\rho^{\prime})$ by virtue of claim (i), just proved.
Since $c^{\prime}=c\in CH(X(D(0,1))$, however, there is a root, $z_{h}$, lying in $D(c^{\prime},\rho^{\prime})$ but not in $D(z_{g},\rho)$, such that $|z_{g}-z_{h}|>\rho\geq\rho^{\prime}-2\delta$ (cf. (21.2)).
Hence $|z_{g}-z_{h}|>3\rho^{\prime}/4$ (implying claim (ii) for the output of Alg. 21.1) because $42\delta\leq\epsilon/4<\rho^{\prime}/$ 4.
Next extend this correctness proof to the general case based on bounds (21.2).
Notice that $c^{\prime}\in D(0,1+\delta)$ because $c\in CH(X(D(0,1)))\subset D(0,1)$ and $|c^{\prime}-c|\leq\delta$.
Hence $\rho=r_{d-m+1}(c^{\prime})\leq 2+\delta$, and so $\rho^{\prime}\leq\rho+2\delta\leq 2+3\delta$ (cf. (21.2)) and $D(c^{\prime},\rho^{\prime})\subseteq D(0,3+4\delta)$.
Therefore $D(c^{\prime},\rho^{\prime})\subseteq D(0,3.5)$ because $8\delta\leq\epsilon\leq 1$ by assumption.
Now recall that #$\#(D(0,3.5))=\#(D(0,1))=m$, while #$\#(D(c^{\prime},\rho^{\prime}))\geq\#(D(c^{\prime},\rho))=m$ for $\rho^{\prime}\geq\rho$ and deduce claim (i) for the output of Alg. 21.1.
Next assume wlog that $c=0$, $c^{\prime}=\delta$, and $z_{g}=c^{\prime}+\rho$ for a root $z_{g}$.
Then $\Re(z_{g})\geq\rho-\delta\geq\rho^{\prime}-3\delta$ because $|c^{\prime}|=|c^{\prime}-c|\leq\delta$ and $\rho^{\prime}\leq\rho+2\delta$ by assumptions.
Furthermore, $\Re(z_{h})<0$ for some root $z_{h}$ because $c=0\in CH(X(D(0,1)))$. (Actually, $\Re(z_{h})$ reaches its maximum $-\frac{z_{h}}{m-1}$ where $z_{h}$ is a simple root and $-\frac{z_{h}}{m-1}$ is a root of multiplicity $m-1$.)
Therefore, $|z_{g}-z_{h}|\geq\Re(z_{g})-\Re(z_{h})>\rho^{\prime}-3\delta\geq\rho^{\prime}-\frac{3}{8}\epsilon$.
This implies claim (ii) for the output of Alg. 21.1 since $\rho^{\prime}>\epsilon$.
21.3 Elaboration upon the compression algorithm
We readily (i) count roots in isolated input disc $D(0,1)$ of Alg. 21.1 by applying the fast root-counter of Sec. 23.4 or alternatively Alg. 7.1 and Prop. 7.1 and (ii) compute a complex point $c^{\prime}$ lying in or near the convex hull of the roots in $C$ by applying fast algorithm of Sec. 23. Finally, we estimate the root radius $\rho^{\prime}:=r_{d-m+1}(c^{\prime})$ by applying Alg. 9.1, which incorporates the $m$-test of of Sec. 23.4, or alternatively Alg. 7.1 and Prop. 7.1.
21.4 Counting $\epsilon$-compression steps
Recall from Sec. 1.3 that the output disc $D(c^{\prime},\rho^{\prime})$ of Alg. 21.1 is said to be an $\epsilon$-disc if $\rho^{\prime}\leq\epsilon$; in this case call the compression by Alg. 21.1 final and output that disc.
Otherwise call the square $S=S(c^{\prime},\rho^{\prime})$ suspect and apply to it subdivision iterations.
Extend Lemma 20.1 to the case where $|z_{g}-z_{h}|\geq\frac{5}{8}\rho^{\prime}$ for a pair of roots $z_{g},z_{h}\in D(c^{\prime},\rho^{\prime})$, combine it with claim (ii) for the output of Alg. 21.1, and obtain that at most $2.5+\log_{2}(8\sigma m/5)$ subdivision steps break the square $S$ into more than one component.
From part (ii) of Sec. 4 recall the bound $22m-$ 2 on the number of compression steps and obtain
@@ 21.4 Counting $\epsilon$-compression steps
For a fixed $\epsilon\geq 0$ let subdivision and compression steps be applied to a component $C$ of suspect squares containing at most $m$ roots. Then the computations involve at most $66m-$ 6 compact components, made up of at most $2424m-$ 24 suspect squares; that is, these computations involve at most $2424m-$ 24 e/i tests.
@@
A compact component stays unbroken and uncompressed in at most $3$ consecutive subdivision steps, which end with a final compression step or follow a non-final compression step and end with breaking a component. Hence all such components are processed in at most $66m-$ 6 subdivision step, each involving at most 4 e/i tests. ∎
22 Counting e/i tests
22.1 An upper estimate
The upper estimate $22m-$ 2 of part (ii) of Sec. 4 on the number of compression steps is sharp – reached in the case of the complete binary tree $\mathbb{T}$. Each compression step is preceded by an e/i test, and so we perform at least $22m-$ 2 e/i steps. According to the following result, which is the main theorem of this section, this lower bound on the number of e/i tests involved is sharp up to a constant factor:
Theorem 22.1.
22.2 The first lemma
We begin our proof of Thm. 22.1 with
Lemma 22.1.
Proof.
[…]
22.3 The second lemma
Lemma 22.2.
Fix two integers $i\geq 0$ and $g\geq 3$, fix a value $\sigma$ of softness, $1<\sigma<\sqrt{2}$, and define the separation fraction $a:=2-\sigma\sqrt{2}$ of Observation 20.1. For all $j$ let $j$-th subdivision step process $v_{j}$ components made up of $w_{j}$ suspect squares, $n_{j}$ of which are non-adjacent, let $\Delta_{j}$ denote their side length, and let the $j$-th steps perform no compression for $j=i,i+1,\dots,i+g$. Then
Proof.
It remains to bound by $O(m)$ the number of e/i tests in the remaining subdivision steps, not interrupted with compression steps.
Furthermore, after a compression step we resume subdivision process with a suspect square whose side length bound decreases in two next subdivision steps from at most $4\Delta_{j}$ to at most $\Delta_{j}$.
22.4 Counting e/i tests at partition steps
Let $h$ denote the height of our partition tree $\mathbb{T}$ and count suspect squares processed in subdivision steps that partition components, temporarily ignoring the other steps.
Notice that
$\sum_{i=0}^{h-g}v_{i+g}\leq\sum_{i=0}^{h}v_{i}-\sum_{i=0}^{g-1}v_{i}<2m-g,$
and so
[…]
Obtain that
Hence
22.5 Extension to counting all e/i steps
Next consider a non-compact component, count all suspect squares processed in all subdivision steps applied until it is partitioned, and then prove in Lemma 22.3 that up to a constant factor the overall number of these suspect squares is bounded by their overall number at all partition steps. Such a result extends to all subdivision steps the bound $n^{\prime}=O(m)$ of (22.3), so far restricted to partition steps. Together with Lemma 22.1 this will complete our proof of Thm. 22.1.
Lemma 22.3.
Let $k+1$ successive subdivision steps with $\sigma$-soft e/i tests be applied to a non-compact component for $\sigma$ of our choice in the range $1<\sigma<\sqrt{2}$. Let the $j$-th step output $n_{j}$ non-adjacent suspect squares for $j=0,1,\dots,k$. Then $\sum_{j=0}^{k-1}n_{j}\leq n_{k}\frac{10\sqrt{2}}{a}$ for the separation fraction $a:=2-\sigma\sqrt{2}$ (cf. Observation 20.1.)
Proof.
Hence $5n_{k}\geq n_{i}a2^{k-i-1.5}$ for all $i$, and so
Remark 22.1.
Remark 22.2.
Alg. 21.1 compresses an input disc $D(0,1)$ into $\eta$-rigid disc for $\eta=5/8$, but our proof of Thm. 22.1 only involves $\eta$ in the argument in Sec. 21.4, and this argument and therefore the whole proof of the theorem can be immediately extended if it relies on compression into any rigid disc.
23 A point near root set and root-count in a disc
We prove that the algorithms of this section work where an input disc is $d^{\nu}$-isolated for a fixed positive constant $\nu$. In application to compression we begin with an isolated disc, say, 3.5-isolated, but can turn it into $d^{\nu}$-isolated for any fixed positive constant $\nu$ by applying $h$ additional subdivision steps or $h$-lifting for $h=O(\log(d))$. This comment can be extended to application of e/i test of Sec. 23.4 instead of Cauchy-based e/i tests of Part II.
23.1 Computation of a point near root set: introductory comments
Recall Schröder’s iterations [132],
$z=y-m~{}{\rm NR}(y),$ (23.1)
which fast converge to a root of multiplicity $m$ [80, Sec. 5.4]. Due to map (1.3) we only need to consider the case where the input disc $D$ is the unit disc $D(0,1)$. We first assume that it contains all $d$ roots and readily verify
Observation 23.1.
Schröder’s single iteration (23.1) for $m=d$ and $p(x)=p_{d}(x-z_{1})^{d}$ outputs the root $z_{1}$ unless $y=z_{1}$.
Next we extend this observation by proving that the output $z$ of Schröder’s iteration (23.1) lies near the convex hull $CH(X(D))$ provided that $m=d$, a disc $D$ contains all $d$ roots, and point $y$ lies far enough from the disc $D$. Here are our specific estimates.
23.2 The case where a disc contains all $d$ roots
Theorem 23.1.
Write $D:=D(0,1)$, $CH:=CH(D)$, and $\Delta:=\Delta(D)$, fix $y$ such that $|y|>2$, let $z:=y-d~{}{\rm NR(y)}$, which is (23.1) for $m=d$. Let $\bar{x}\in CH$. Then $|z-\bar{x}|\leq\alpha~{}\Delta$ for $\alpha:=\frac{(|y|-1)^{2}}{(y|-2)^{2}}=(1+\frac{1}{(y|-2})^{2}$, and so $\alpha=2.25$ for $u|=4$ and $\alpha=25/16$ for $u|=6$, $\alpha\mapsto 1$ $|y|\mapsto\infty$.
Proof.
Recall that $|\bar{x},z_{1},\dots,z_{d}\in D(0,1)$, $|z_{j}|\leq 1$ and $|\bar{x}-z_{j}|\leq\Delta$ for $j=1,\dots,d$. Hence (cf. (3.1))
@@ 23.2 The case where a disc contains all $d$ roots
Under the assumptions of Thm. 23.2 the disc $D(z,r_{1}(z))$ is $\frac{1}{2+2\alpha}$-rigid.
@@
Now we devise our fast compression algorithm inthe case where $m=d$.

### v37

@@ 15.3 The bit operation complexity estimates
In this part of the paper we decrease the number of subdivision steps of our root-finders of Parts II and III by a factor of $b/\log(b)$, thus arriving at our Main Theorems 1 – 4.
In Sec. 17 we extend the outline of this acceleration from Sec. 4. It relies on an algorithm for compression of an isolated disc, which we devise and analyze also in Sec. 17. The algorithm involves estimation of the $\ell$th smallest root radius, which is of independent interest and is covered in Sec. 16. It is reduced to recursive application of an $\ell$-test; we present it in Sec. 13 for any positive $\ell\leq d$ and any input disc, but in Sec. 18 we specify simpler tests that are sufficient in application to compression. In Sec. 20 we estimate the overall complexity of the resulting root-finder, which intertwines at most $22m-$ 2 compression steps with e/i tests, whose number we estimate in Sec. 19. In Sec. 21 we further accelerate polynomial root-finding for a large class of inputs.
16 Estimation of the $\ell$-th smallest root radius
In this section we narrow a fixed range $[0,R]$ for the $\ell$-th smallest root radius $r_{d-\ell+1}(c)$ from a fixed complex point $c$. To simplify our exposition we let $R=1$ and $c=0$, wlog due to (1.4).
Problem RR. Compute a range $[\rho_{-},\rho_{+}]$ such that $0\leq\rho_{-}\leq r_{d-\ell+1}\leq\rho_{+}$ and $\rho_{+}\leq\epsilon=1/2^{b}$ or $\rho_{+}\leq\beta\rho_{-}$ for fixed $b>0$ and $\beta=2^{\phi}>1$, provided that we are given a positive integer $\ell\leq d$, initial range $[0,1]$ for $r_{d-\ell+1}$, and a $\sigma$-soft $\ell$-test for a fixed $\sigma>1$.
We first use this test to check whether $r_{d-\ell+1}\leq 1/2^{b}$. If so, we are done. Otherwise we seek $r_{d-\ell+1}$ in the range $[\frac{1}{2^{b}\sigma},1]$. We apply binary search, said to be bisection of logarithms or BoL, to the range $[-\bar{b},0]=[-\log_{2}(\frac{\sigma}{\epsilon}),\log_{2}(1)]$, for $\bar{b}=\log_{2}(\frac{\sigma}{\epsilon})$, to approximate $\log_{2}(r_{d-\ell+1})$.
Since the test is soft, so is BoL as well: for a range $[b_{-},b_{+}]$ with midpoint $\widehat{b}=0.5(b_{+}+b_{-})$ regular rigid bisection continues the search in the range $[b_{-},\widehat{b}]$ or $[\widehat{b},b_{+}]$, but our soft BoL continues it in the range $[b_{-},(1+\alpha)\widehat{b}]$ or $[(1-\alpha)\widehat{b},b_{+}]$ for a fixed small positive $\alpha$. Accordingly we slightly increase the range of search for $r_{d-\ell+1}$ to $[-\bar{b},0]$ for $\bar{b}:=\log_{2}(\frac{\sigma}{\epsilon})$. BoL still converges linearly and just a little slower than the regular rigid bisection: the approximation error decreases by a factor of $\frac{2}{1+\alpha}$ rather than by twice. [^27] If we applied BoL to the range $[0,1]$, we would have to bisect the infinite range $[-\infty,0]$, but we deal with the narrower range $[-b,0]$ since we have narrowed the range $[0,1]$ to $[\epsilon,1]$.
Algorithm 16.1.
Bracketing the $\ell$-th smallest root radius based on $\ell$-test.
INPUT: An integer $\ell$ such that $1\leq\ell\leq d$ and #$\#(D(0,1))\leq\ell$, $\sigma=1+\alpha$ for $20<\alpha<1/$ 2, a $\sigma$-soft $\ell$-test0,ρ for a positive $\rho$ (see Sec. 13.4 for such tests), and two tolerance values $\beta:=2^{\phi}>1$ and $\epsilon=1/2^{b}$, $b\geq 1$.
OUTPUT: A solution $[\rho_{-},\rho_{+}]$ of Problem RR.
[…]
The disc $D(0,\epsilon)$ contains at least $\ell$ roots, that is, the range $[0,\epsilon]$ brackets the $\ell$th smallest root radius $r_{d-\ell+1}$ if the $\sigma$-soft $\ell$-test applied to the disc $D(0,\frac{\epsilon}{\sigma})$ outputs 1. Otherwise, the disc $D(0,\frac{\epsilon}{\sigma})$ contains less than $\ell$ roots, and so the root radius $r_{d-\ell+1}$ lies in the range $[\frac{\epsilon}{\sigma},1]$, in which case we continue our search.
Readily verify that our policy of recursively updating the range $[b_{-},b_{+}]$ maintains this property throughout and that updating stops where $b_{+}-b_{-}\leq\phi$, that is, where $\rho_{+}\leq\beta\rho_{-}$. This implies correctness of Alg. 16.1.
Let us estimate the number $\mathbb{I}$ of the invocations of an $\ell$-test.
Proposition 16.1.
Alg. 16.1 invokes an $\ell$-test $\mathbb{I}$ times where $\mathbb{I}\leq\lceil\log_{\nu}(\frac{\bar{b}}{\phi})\rceil+1$ for $\nu=\frac{2}{1+\alpha}$, and so $\mathbb{I}=O(\log(b))$ provided that $\nu-1$, $\sigma-1$, and $\phi$ are positive constants.
Proof.
Alg. 16.1 stops if the $\ell$-test outputs 1. Otherwise the test decreases the length $\bar{b}=b_{+}-b_{-}$ of its input range $[b_{-},b_{+}]$ for the exponent to $0.5(1+\alpha)\bar{b}$. $i$ applications change it into $\frac{(1+\alpha)^{i}}{2^{i}}\bar{b}$ for $i=1,2,\dots$, and the proposition follows. ∎
Remark 16.1.
Clearly, $\mathbb{I}$ applications of $\ell$-test increase the number of evaluation points for the ratio $\frac{p^{\prime}(x)}{p(x)}$ by at most a factor of $\mathbb{I}$, and if the $\ell$-test is randomized, then the error probability also grows by at most a factor of $\mathbb{I}$. Let us specify these estimates for Alg. 16.1.
Corollary 16.1.
Let precisely $\ell$ zeros of a black box polynomial $p(x)$ of a degree $d$ lie in the unit disc, that is, let #$\#(D(0,1))=\ell$, and let $i(D(0,1))\geq\sigma$, for $\sigma-1$ exceeding a positive constant. Given two positive tolerance values $\epsilon=1/2^{b}$ and $\phi$ and a black box $\sigma$-soft $\ell$-test, which evaluates the polynomial $p(x)$ at $\mathcal{NR}$ points, Alg. 16.1 approximates the $\ell$-th smallest root radius $|z_{j}|$ within the relative error $2^{\phi}$ or determines that $|z_{j}|\leq\epsilon$, at the cost of evaluation of polynomial $p(x)$ at $\mathbb{I}\cdot\mathcal{NR}$ points for $\mathbb{I}$ estimated in Prop. 16.1, $\mathbb{I}=O(\log(b))$ provided that $\nu-1$ and $\phi$ are positive constants.
17 Compression of an isolated disc
17.1 A motivation: compact components in the union of suspect squares
As in Sec. 4, partition the union of suspect squares at every subdivision iteration into components and associate breaking (partition of) components into sub-components with component tree, which has at most $22m-$ 2 edges. Next we will bound the numbers of subdivision steps (and hence of e/i tests) at and between partitions of components.
@@ 17.1 A motivation: compact components in the union of suspect squares
We call a pair of suspect squares non-adjacent if they share no vertices. Call a component compact if it contains just a single non-adjacent suspect square, that is, consists of at most four suspect squares sharing a vertex (see Fig. 5).
@@
A component of suspect squares is broken in at most $1+\lfloor 0.5+\log_{2}(m)\rfloor$ successive subdivision steps unless it is compact.
Observation 17.2.
@@
Cor. 17.1 bounds the number of subdivision steps between partitions of non-compact components, but compact components can be processed very slowly (see Fig. 3). For a remedy, we intertwine subdivision process with compression of MCD of a component that stays compact in three subdivision steps, in which case the MCD is $4\sqrt{2}$-isolated by virtue of Observation 17.2.
Next we elaborate upon such compression of MCDs.
17.2 The compression problem and flowchart of its solution
Problem COMPR. INPUT: a positive $\epsilon\leq 1$, a constant $\eta>0$, a black box polynomial $p(x)$ of a degree $d$, and an isolated disc $D$ with a non-empty root set.
@@ 17.2 The compression problem and flowchart of its solution
OUTPUT: an integer $m$, a complex $c\in D$, and positive $\rho$ and $\eta$ such that #$\#(D(c,\rho))=\#(D)$ and the disc $D(c,\rho)$ is $\eta$-rigid unless $\rho\leq\epsilon$.
Flowchart of a compression algorithm. To solve Problem COMPR we successively compute (a) $m=\#(D)$, (b) complex point $c$ lying in or near the minimal covering disc of the root set $X(D)$, and (c) positive $\rho$ and and $\eta$ such that #$\#(D(c,\rho))=m$ and the disc $D(c,\rho)$ is $\eta$-rigid unless $\rho\leq\epsilon$.
Wlog assume that $D$ is a unit disc $D(0,1)$ and (a) compute $m=\#(D)$ for isolated unit disc $D:=D(0,1)$ by using $O(\log(d))$ queries of an evaluation oracle (see Observations 12.1 and 12.2).
Next we elaborate upon the algorithms for stages (b) and (c).
17.3 A point in or near the minimal covering disc for the $d$ roots
Definition 17.1.
A disc is said to be strongly isolated if it is $d^{\nu}$-isolated for a sufficiently large constant $\nu$ of our choice. In particular every disc containing all $d$ roots is strongly isolated.
Next we successively compress the unit disc $D=D(0,1)$ in the s where it contains $d$ roots, is strongly isolated, and is just isolated.
Schröder’s iterations [104],
$z=y-m~{}{\rm NR}(y),$ (17.1)
fast converge to a root of multiplicity $m$ [63, Sec. 5.4] and for $m=d$ outputs a point in or near the MSD of the the set of the $d$ roots if point $y$ lies far enough from that set:
@@ 17.3 A point in or near the minimal covering disc for the $d$ roots
Write $D:=D(0,1)$ and $\Delta:=\Delta(X(D))$, let #$\#(D)=d$, fix $y$ such that $|y|>2$, let $z:=y-d~{}{\rm NR(y)}$ and $\bar{x}\in MCD(D)$. Then $|z-\bar{x}|\leq\alpha~{}\Delta$ for $\alpha:=\frac{(|y|-1)^{2}}{(y|-2)^{2}}=(1+\frac{1}{(y|-2})^{2}$, and so $\alpha=2.25$ for $|y|=4$, $\alpha=25/16$ for $|y|=6$, and $\alpha\mapsto 1$ as $|y|\mapsto\infty$.
@@
Recall that $\bar{x},z_{1},\dots,z_{d}\in X(D)$, $|z_{j}|\leq 1$ and $|\bar{x}-z_{j}|\leq\Delta$ for $j=1,\dots,d$. Hence (cf. (3.1))
@@
Under the assumptions of Thm. 17.2 the disc $D(z,r_{1}(z))$ is $\frac{1}{2+2\alpha}$-rigid.
@@
17.4 Compression algorithm for a disc containing $d$ roots
Next we compress a disc containing $d$ roots towards their minimal covering disc.